# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 6.2.1 Block Header and Block Body
  • Block Header: The block header contains metadata about the block[144]. It includes:
@@
  • Block Body: The block body contains the actual data, typically a list of validated transactions[144]. Each transaction records details like the sender, receiver, and the amount transferred.
[…]
@@
The immutability of the blockchain arises from the fact that each block is cryptographically linked to the previous block through its hash[145]. If any data within a block is tampered with, the hash of that block will change, breaking the link with the subsequent block. This cascading effect requires recalculating the hashes of all subsequent blocks, which is computationally infeasible in most blockchain networks, particularly those using proof-of-work or similar consensus mechanisms. Therefore, the structure of the blockchain ensures that once data is added to the chain, it cannot be altered without redoing the entire chain, making it effectively tamper-resistant.
@@ 6.2.2 Block Hash and Previous Block Hash
Every block has a unique identifier, known as the block hash[146]. This hash is created by applying a cryptographic function (such as SHA-256) to the block’s contents, including its header and body.
@@ 6.3.1 Structure of the Merkle Tree
A Merkle tree (also known as a binary hash tree) is a data structure used in blockchain to efficiently and securely verify the integrity of transactions[147]. Each leaf node of the Merkle tree contains a hash of a single transaction, while each non-leaf node contains a hash of its two child nodes.
@@ 6.3.3 Application of Merkle Tree in Data Verification
Merkle trees enable efficient data verification without requiring the entire dataset[148]. In the context of blockchain, a Merkle tree is used to verify if a specific transaction is part of a block without needing to download the entire block. This is particularly useful for light nodes, which only store the block headers rather than the full blockchain data.
@@ 6.4.1 Definition and Purpose of PoW
Proof of Work (PoW) is a consensus algorithm used in blockchain networks, such as Bitcoin, to ensure network security and verify transactions[149]. It requires miners (nodes that validate transactions) to solve a computational puzzle, where the goal is to find a value that, when hashed, meets a difficulty target set by the network.
@@ 6.4.2 Nonce Selection and Hash Calculation
The key element of PoW is finding a nonce—a random number that miners adjust to solve the computational puzzle[150]. The task is to find a nonce such that when the block header (including the nonce) is hashed, the resulting hash is below a certain threshold, known as the difficulty target.
@@
7difficulty_target = "$0$ffffffffffffffffffffffffffffffffffffffffffffffffffffffffffff"
@@ 6.5.1 Timestamp in Blockchain
In a blockchain, each block contains a timestamp that indicates when the block was created. The timestamp plays a crucial role in ordering blocks and transactions chronologically. This is essential because the timestamp ensures that the chain of blocks maintains a proper sequence and prevents issues like replay attacks, where older transactions could be reused maliciously[126].
The timestamp is included in the block header, along with other data such as the previous block hash, the Merkle root, and the nonce[151]. Here’s an example of a block header structure:
@@ 6.5.2 Immutability of Blockchain
One of the fundamental principles of blockchain is immutability, meaning that once data is written to a block and added to the blockchain, it cannot be altered or deleted without invalidating the entire chain. This is achieved through the use of cryptographic hashes and the PoW consensus mechanism.[152]
@@ 7.1.1 Definition of Inputs and Outputs
A Bitcoin transaction is a transfer of value between Bitcoin addresses. It consists of two main components: inputs and outputs. The input defines the source of the funds, and the output defines where the funds are being sent. More precisely, inputs refer to references to previous unspent outputs (UTXOs - Unspent Transaction Outputs), and outputs specify the destination addresses and the amount of Bitcoin being transferred[153].
Inputs: Each input in a Bitcoin transaction points to an unspent output from a previous transaction[154]. This input consumes the value of that output to be spent in the current transaction. An input also includes a cryptographic signature, which proves ownership of the funds associated with the output.
@@ 7.1.2 Explanation of the UTXO Model
The Bitcoin network relies on the UTXO (Unspent Transaction Output) model to track the ownership of Bitcoin[156]. Unlike traditional accounts where balances are directly managed, Bitcoin transactions use UTXOs to ensure accuracy, security, and the prevention of double-spending.
@@ 7.2.1 Role of Full Nodes
A full node is a computer that fully validates transactions and blocks in the Bitcoin network[157]. Full nodes download and store the entire blockchain and participate in transaction validation, ensuring that the rules of the Bitcoin protocol are strictly followed.
@@ 7.2.2 Light Nodes and Simplified Payment Verification (SPV)
Light nodes, also known as SPV (Simplified Payment Verification) nodes[158], are a type of node that does not store the full blockchain but instead relies on full nodes for transaction verification.
Light nodes only download the block headers (which are much smaller in size) instead of the entire blockchain[159]. This allows them to verify transactions without needing to download and store all the data. When a light node wants to verify a transaction, it can request proof from full nodes using Merkle trees, a cryptographic structure that ensures the integrity of the data.
@@ 7.3.1 Transaction Propagation in the P2P Network
Bitcoin uses a peer-to-peer (P2P) network to propagate transactions and blocks[129]. When a user broadcasts a transaction, it is first sent to the nearest nodes in the network. Each node then verifies the transaction and relays it to its peers. This process continues until the transaction reaches miners, who will include it in the next block.
@@ 7.3.3 Six Confirmations Rule in Blockchain
In the Bitcoin network, there is a commonly accepted practice known as the six confirmations rule[160]. This rule states that after a transaction has been confirmed by six blocks, it is considered extremely secure and unlikely to be reversed. The reasoning behind this rule is based on the probability of an attacker successfully rewriting the blockchain. With each added block, the chances of an attacker succeeding diminish, making six confirmations a highly secure threshold.
@@ 7.4.1 Mechanism of Mining
Bitcoin mining is the process by which new transactions are added to the Bitcoin blockchain. It involves miners competing to solve a complex cryptographic puzzle based on the Proof of Work (PoW) consensus mechanism[161]. The puzzle requires miners to find a hash that, when applied to the block’s header, meets certain conditions (i.e., it must be lower than a certain target). This process is computationally intensive and ensures that adding new blocks to the blockchain requires substantial effort.
@@
The difficulty adjustment mechanism is crucial to Bitcoin’s stability[162]. It ensures that blocks are added to the blockchain approximately every 10 minutes, regardless of the total computational power in the network. Every 2016 blocks (approximately every two weeks), the difficulty is adjusted based on how quickly or slowly the previous blocks were mined. If the blocks were mined too quickly, the difficulty increases; if too slowly, the difficulty decreases. This self-regulating process ensures a stable and predictable issuance of new bitcoins.
@@ 7.4.2 Miner Rewards and Block Rewards
  • Block Subsidy: The block subsidy is a fixed amount of newly minted bitcoins. Initially set at 50 bitcoins per block, the subsidy is halved approximately every four years during an event called the halving[163]. After the third halving in 2020, the block subsidy is now 6.25 bitcoins per block. The subsidy will continue to halve every 210,000 blocks until it eventually reaches zero, around the year 2140.
@@ 7.4.3 Bitcoin’s Total Supply and Scarcity
Bitcoin’s economic model is built on the principle of scarcity[164]. There will only ever be 21 million bitcoins in existence. This limit is hardcoded into the Bitcoin protocol and cannot be changed without a complete overhaul of the system, which would require consensus from the entire network.
@@ 7.5.1 Generation of Private Keys, Public Keys, and Addresses
A Bitcoin wallet is essentially a pair of cryptographic keys: a private key and a public key[165]. These keys are generated through the use of cryptographic algorithms, primarily elliptic curve cryptography (ECC).
@@
  • Cold Wallets: Cold wallets are offline storage solutions, such as hardware wallets or paper wallets. Because they are not connected to the internet, they are immune to hacking attacks and are generally considered more secure[166]. However, they are less convenient for frequent transactions. An example of a cold wallet is a hardware wallet like Trezor[137] or Ledger[136].
  • Hot Wallets: Hot wallets are online wallets connected to the internet, such as software wallets or exchange-based wallets. While they provide easy access for frequent transactions, they are more vulnerable to security risks like hacking. Popular examples include mobile apps like Trust Wallet or web-based wallets provided by exchanges like Binance[167] or Coinbase[168].
@@ 7.5.3 Multi-Signature Wallets
Multi-signature (multisig) wallets provide an additional layer of security by requiring multiple private keys to authorize a transaction[169]. In a traditional Bitcoin wallet, a single private key is used to sign transactions. In contrast, a multisig wallet requires approval from multiple parties[170].
@@ 8.1.1 Satoshi Nakamoto’s Genesis Message
One of the most remarkable features of Bitcoin’s origin is the message embedded in the first block, known as the Genesis Block or Block 0[146], by the pseudonymous creator Satoshi Nakamoto. The Genesis Block, mined on January 3, 2009, contains a hidden message in its coinbase parameter:
@@
This message is a reference to the headline of an article published in the UK newspaper *The Times* on the same day[171]. It reflects the financial instability during the 2008 global financial crisis, where many banks were bailed out by governments in response to their near-collapse.
The message carries symbolic significance, representing Bitcoin’s foundational goal: to create an alternative financial system that is decentralized, transparent, and free from the control of central banks and governments[172]. By embedding this message, Satoshi Nakamoto highlighted Bitcoin’s intent to be a solution to the failings of the traditional financial system, which had shown vulnerabilities during the crisis.
@@ 8.2.1 Block Size Limit and Transaction Speed
Bitcoin’s original design includes a 1MB limit on the size of each block[173], meaning that only a certain number of transactions can fit into a block. On average, Bitcoin processes around 7 transactions per second (TPS)[174]. During times of low network activity, this transaction rate is adequate. However, during periods of high demand, the limited block size causes a bottleneck, leading to slower transaction speeds and higher fees.
@@
Bitcoin’s block size constraint directly impacts its ability to scale[175]. When more transactions are submitted than the network can process, users experience longer waiting times for their transactions to be confirmed. This issue became particularly noticeable during the 2017 bull market, when transaction fees skyrocketed due to network congestion.
@@ 8.2.2 Technical Challenges of Bitcoin Scalability
  • Trade-off between decentralization and scalability: As Bitcoin’s block size is increased, more data is required to be processed and stored by full nodes. This can lead to centralization, as only those with significant computing power and storage capabilities can afford to run a full node. In contrast, a smaller block size allows more users to participate in the network, maintaining decentralization but at the cost of slower transaction speedse[176].
  • Security risks: A key aspect of Bitcoin’s security is its Proof of Work (PoW) consensus mechanism. Increasing the block size or changing how transactions are processed could introduce new vulnerabilities. For example, larger blocks might take longer to propagate across the network, leading to temporary forks and security riskse[177].
@@
    • Segregated Witness (SegWit): SegWite[178] was introduced to increase the effective block size without increasing the actual block size limit. It achieves this by removing certain parts of the transaction data, making more space for transactions in each block.
    • The Lightning Network: The Lightning Networke[179] is a second-layer solution that allows for off-chain transactions. By creating payment channels between users, the Lightning Network can handle a large number of transactions outside the blockchain, only settling on-chain when necessary. This significantly reduces the burden on the Bitcoin network, allowing for faster and cheaper transactions.
    • Hard forks (e.g., Bitcoin Cash): In 2017, a group of developers initiated a hard fork of Bitcoin, resulting in the creation of Bitcoin Cash (BCH)e[180]. Bitcoin Cash increased the block size to 8MB and later to 32MB, allowing for more transactions per block. However, this came with its own set of challenges, including reduced decentralization and potential security risks.
@@ 8.3.1 Definition and Examples of Soft Forks
Soft forks are backward-compatible upgrades to a blockchain protocole[181]. This means that non-upgraded nodes (those that have not adopted the changes) can still interact with upgraded nodes without breaking the network’s overall functionality. Soft forks typically introduce new rules that are more restrictive than before, allowing older nodes to continue functioning as long as they follow the new rules.
For example, the BIP66 soft fork was implemented in Bitcoin in 2015[182]. It introduced strict DER (Distinguished Encoding Rules) signatures, which enforced stricter rules for the way digital signatures were encoded on the Bitcoin network. Before BIP66, there was a degree of flexibility in how these signatures were generated. This flexibility sometimes caused issues, such as minor differences in how signatures were processed, leading to bugs and potential vulnerabilities. By enforcing a strict standard, BIP66 helped prevent these issues while remaining compatible with non-upgraded nodes.
Another notable soft fork was the SegWit (Segregated Witness)[183] upgrade, which will be explained in more detail in a later subsection.

### v3

@@ 6.2.1 Block Header and Block Body
  • Block Header: The block header contains metadata about the block[Gupta2021BlockchainTP]. It includes:
@@
  • Block Body: The block body contains the actual data, typically a list of validated transactions[Gupta2021BlockchainTP]. Each transaction records details like the sender, receiver, and the amount transferred.
Block HeaderBlock BodyHash: $H_{1}$Block HeaderBlock BodyHash: $H_{2}$Block HeaderBlock BodyHash: $H_{3}$$H_{1}$$H_{2}$
@@
The immutability of the blockchain arises from the fact that each block is cryptographically linked to the previous block through its hash[Kizza2020BlockchainsCA]. If any data within a block is tampered with, the hash of that block will change, breaking the link with the subsequent block. This cascading effect requires recalculating the hashes of all subsequent blocks, which is computationally infeasible in most blockchain networks, particularly those using proof-of-work or similar consensus mechanisms. Therefore, the structure of the blockchain ensures that once data is added to the chain, it cannot be altered without redoing the entire chain, making it effectively tamper-resistant.
@@ 6.2.2 Block Hash and Previous Block Hash
Every block has a unique identifier, known as the block hash[C2018AnOO]. This hash is created by applying a cryptographic function (such as SHA-256) to the block’s contents, including its header and body.
@@ 6.3.1 Structure of the Merkle Tree
A Merkle tree (also known as a binary hash tree) is a data structure used in blockchain to efficiently and securely verify the integrity of transactions[Liu2021MerkleTA]. Each leaf node of the Merkle tree contains a hash of a single transaction, while each non-leaf node contains a hash of its two child nodes.
@@ 6.3.3 Application of Merkle Tree in Data Verification
Merkle trees enable efficient data verification without requiring the entire dataset[Wang2023ATD]. In the context of blockchain, a Merkle tree is used to verify if a specific transaction is part of a block without needing to download the entire block. This is particularly useful for light nodes, which only store the block headers rather than the full blockchain data.
@@ 6.4.1 Definition and Purpose of PoW
Proof of Work (PoW) is a consensus algorithm used in blockchain networks, such as Bitcoin, to ensure network security and verify transactions[ferdous2020blockchainconsensusalgorithmssurvey]. It requires miners (nodes that validate transactions) to solve a computational puzzle, where the goal is to find a value that, when hashed, meets a difficulty target set by the network.
@@ 6.4.2 Nonce Selection and Hash Calculation
The key element of PoW is finding a nonce—a random number that miners adjust to solve the computational puzzle[Raza2021EnergyEM]. The task is to find a nonce such that when the block header (including the nonce) is hashed, the resulting hash is below a certain threshold, known as the difficulty target.
@@
7difficulty_target = "0000ffffffffffffffffffffffffffffffffffffffffffffffffffffffffffff"
@@ 6.5.1 Timestamp in Blockchain
In a blockchain, each block contains a timestamp that indicates when the block was created. The timestamp plays a crucial role in ordering blocks and transactions chronologically. This is essential because the timestamp ensures that the chain of blocks maintains a proper sequence and prevents issues like replay attacks, where older transactions could be reused maliciously[Hunt2008BitcoinAP].
The timestamp is included in the block header, along with other data such as the previous block hash, the Merkle root, and the nonce[Regnath2020BlockchainWT]. Here’s an example of a block header structure:
@@ 6.5.2 Immutability of Blockchain
One of the fundamental principles of blockchain is immutability, meaning that once data is written to a block and added to the blockchain, it cannot be altered or deleted without invalidating the entire chain. This is achieved through the use of cryptographic hashes and the PoW consensus mechanism.[Rahman2020DesignPF]
@@ 7.1.1 Definition of Inputs and Outputs
A Bitcoin transaction is a transfer of value between Bitcoin addresses. It consists of two main components: inputs and outputs. The input defines the source of the funds, and the output defines where the funds are being sent. More precisely, inputs refer to references to previous unspent outputs (UTXOs - Unspent Transaction Outputs), and outputs specify the destination addresses and the amount of Bitcoin being transferred[bitcoinDoesBitcoin].
Inputs: Each input in a Bitcoin transaction points to an unspent output from a previous transaction[Kang2023BitcoinDA]. This input consumes the value of that output to be spent in the current transaction. An input also includes a cryptographic signature, which proves ownership of the funds associated with the output.
@@ 7.1.2 Explanation of the UTXO Model
The Bitcoin network relies on the UTXO (Unspent Transaction Output) model to track the ownership of Bitcoin[DelgadoSegura2018AnalysisOT]. Unlike traditional accounts where balances are directly managed, Bitcoin transactions use UTXOs to ensure accuracy, security, and the prevention of double-spending.
@@ 7.2.1 Role of Full Nodes
A full node is a computer that fully validates transactions and blocks in the Bitcoin network[Abe2018MitigatingBN]. Full nodes download and store the entire blockchain and participate in transaction validation, ensuring that the rules of the Bitcoin protocol are strictly followed.
@@ 7.2.2 Light Nodes and Simplified Payment Verification (SPV)
Light nodes, also known as SPV (Simplified Payment Verification) nodes[AlBassam2018FraudAD], are a type of node that does not store the full blockchain but instead relies on full nodes for transaction verification.
Light nodes only download the block headers (which are much smaller in size) instead of the entire blockchain[Zamyatin2020TxChainEC]. This allows them to verify transactions without needing to download and store all the data. When a light node wants to verify a transaction, it can request proof from full nodes using Merkle trees, a cryptographic structure that ensures the integrity of the data.
@@ 7.3.1 Transaction Propagation in the P2P Network
Bitcoin uses a peer-to-peer (P2P) network to propagate transactions and blocks[Donet2014TheBP]. When a user broadcasts a transaction, it is first sent to the nearest nodes in the network. Each node then verifies the transaction and relays it to its peers. This process continues until the transaction reaches miners, who will include it in the next block.
@@ 7.3.3 Six Confirmations Rule in Blockchain
In the Bitcoin network, there is a commonly accepted practice known as the six confirmations rule[Wan2019EvaluatingTI]. This rule states that after a transaction has been confirmed by six blocks, it is considered extremely secure and unlikely to be reversed. The reasoning behind this rule is based on the probability of an attacker successfully rewriting the blockchain. With each added block, the chances of an attacker succeeding diminish, making six confirmations a highly secure threshold.
@@ 7.4.1 Mechanism of Mining
Bitcoin mining is the process by which new transactions are added to the Bitcoin blockchain. It involves miners competing to solve a complex cryptographic puzzle based on the Proof of Work (PoW) consensus mechanism[Ankalkoti2017ARS]. The puzzle requires miners to find a hash that, when applied to the block’s header, meets certain conditions (i.e., it must be lower than a certain target). This process is computationally intensive and ensures that adding new blocks to the blockchain requires substantial effort.
@@
The difficulty adjustment mechanism is crucial to Bitcoin’s stability[Zhang2018AnAO]. It ensures that blocks are added to the blockchain approximately every 10 minutes, regardless of the total computational power in the network. Every 2016 blocks (approximately every two weeks), the difficulty is adjusted based on how quickly or slowly the previous blocks were mined. If the blocks were mined too quickly, the difficulty increases; if too slowly, the difficulty decreases. This self-regulating process ensures a stable and predictable issuance of new bitcoins.
@@ 7.4.2 Miner Rewards and Block Rewards
  • Block Subsidy: The block subsidy is a fixed amount of newly minted bitcoins. Initially set at 50 bitcoins per block, the subsidy is halved approximately every four years during an event called the halving[Meynkhard2019FairMV]. After the third halving in 2020, the block subsidy is now 6.25 bitcoins per block. The subsidy will continue to halve every 210,000 blocks until it eventually reaches zero, around the year 2140.
@@ 7.4.3 Bitcoin’s Total Supply and Scarcity
Bitcoin’s economic model is built on the principle of scarcity[Sedliaik2023FactorsSB]. There will only ever be 21 million bitcoins in existence. This limit is hardcoded into the Bitcoin protocol and cannot be changed without a complete overhaul of the system, which would require consensus from the entire network.
@@ 7.5.1 Generation of Private Keys, Public Keys, and Addresses
A Bitcoin wallet is essentially a pair of cryptographic keys: a private key and a public key[Gutoski2015HierarchicalDB]. These keys are generated through the use of cryptographic algorithms, primarily elliptic curve cryptography (ECC).
@@
  • Cold Wallets: Cold wallets are offline storage solutions, such as hardware wallets or paper wallets. Because they are not connected to the internet, they are immune to hacking attacks and are generally considered more secure[Barakat2022COMPARISONOH]. However, they are less convenient for frequent transactions. An example of a cold wallet is a hardware wallet like Trezor[trezorTrezorHardware] or Ledger[ledgerHardwareWallet].
  • Hot Wallets: Hot wallets are online wallets connected to the internet, such as software wallets or exchange-based wallets. While they provide easy access for frequent transactions, they are more vulnerable to security risks like hacking. Popular examples include mobile apps like Trust Wallet or web-based wallets provided by exchanges like Binance[binance] or Coinbase[coinbase].
@@ 7.5.3 Multi-Signature Wallets
Multi-signature (multisig) wallets provide an additional layer of security by requiring multiple private keys to authorize a transaction[Goel2023MultisignatureCW]. In a traditional Bitcoin wallet, a single private key is used to sign transactions. In contrast, a multisig wallet requires approval from multiple parties[Tran2023DeploymentOT].
@@ 8.1.1 Satoshi Nakamoto’s Genesis Message
One of the most remarkable features of Bitcoin’s origin is the message embedded in the first block, known as the Genesis Block or Block 0[C2018AnOO], by the pseudonymous creator Satoshi Nakamoto. The Genesis Block, mined on January 3, 2009, contains a hidden message in its coinbase parameter:
@@
This message is a reference to the headline of an article published in the UK newspaper *The Times* on the same day[raskin2018digital]. It reflects the financial instability during the 2008 global financial crisis, where many banks were bailed out by governments in response to their near-collapse.
The message carries symbolic significance, representing Bitcoin’s foundational goal: to create an alternative financial system that is decentralized, transparent, and free from the control of central banks and governments[Maurer2015perhapsTR]. By embedding this message, Satoshi Nakamoto highlighted Bitcoin’s intent to be a solution to the failings of the traditional financial system, which had shown vulnerabilities during the crisis.
@@ 8.2.1 Block Size Limit and Transaction Speed
Bitcoin’s original design includes a 1MB limit on the size of each block[GroupBlockSI], meaning that only a certain number of transactions can fit into a block. On average, Bitcoin processes around 7 transactions per second (TPS)[Gbel2017IncreasedBS]. During times of low network activity, this transaction rate is adequate. However, during periods of high demand, the limited block size causes a bottleneck, leading to slower transaction speeds and higher fees.
@@
Bitcoin’s block size constraint directly impacts its ability to scale[Croman2016OnSD]. When more transactions are submitted than the network can process, users experience longer waiting times for their transactions to be confirmed. This issue became particularly noticeable during the 2017 bull market, when transaction fees skyrocketed due to network congestion.
@@ 8.2.2 Technical Challenges of Bitcoin Scalability
  • Trade-off between decentralization and scalability: As Bitcoin’s block size is increased, more data is required to be processed and stored by full nodes. This can lead to centralization, as only those with significant computing power and storage capabilities can afford to run a full node. In contrast, a smaller block size allows more users to participate in the network, maintaining decentralization but at the cost of slower transaction speedse[Chohan2019TheLT].
  • Security risks: A key aspect of Bitcoin’s security is its Proof of Work (PoW) consensus mechanism. Increasing the block size or changing how transactions are processed could introduce new vulnerabilities. For example, larger blocks might take longer to propagate across the network, leading to temporary forks and security riskse[Lewenberg2015InclusiveBC].
@@
    • Segregated Witness (SegWit): SegWite[Kedziora2022AnalysisOS] was introduced to increase the effective block size without increasing the actual block size limit. It achieves this by removing certain parts of the transaction data, making more space for transactions in each block.
    • The Lightning Network: The Lightning Networke[Bartolucci2019APM] is a second-layer solution that allows for off-chain transactions. By creating payment channels between users, the Lightning Network can handle a large number of transactions outside the blockchain, only settling on-chain when necessary. This significantly reduces the burden on the Bitcoin network, allowing for faster and cheaper transactions.
    • Hard forks (e.g., Bitcoin Cash): In 2017, a group of developers initiated a hard fork of Bitcoin, resulting in the creation of Bitcoin Cash (BCH)e[Aggarwal2019ASA]. Bitcoin Cash increased the block size to 8MB and later to 32MB, allowing for more transactions per block. However, this came with its own set of challenges, including reduced decentralization and potential security risks.
@@ 8.3.1 Definition and Examples of Soft Forks
Soft forks are backward-compatible upgrades to a blockchain protocole[Zamyatin2018AWV]. This means that non-upgraded nodes (those that have not adopted the changes) can still interact with upgraded nodes without breaking the network’s overall functionality. Soft forks typically introduce new rules that are more restrictive than before, allowing older nodes to continue functioning as long as they follow the new rules.
For example, the BIP66 soft fork was implemented in Bitcoin in 2015[bitcoinSoftforkBitcoin]. It introduced strict DER (Distinguished Encoding Rules) signatures, which enforced stricter rules for the way digital signatures were encoded on the Bitcoin network. Before BIP66, there was a degree of flexibility in how these signatures were generated. This flexibility sometimes caused issues, such as minor differences in how signatures were processed, leading to bugs and potential vulnerabilities. By enforcing a strict standard, BIP66 helped prevent these issues while remaining compatible with non-upgraded nodes.
Another notable soft fork was the SegWit (Segregated Witness)[bitcoin0141Bitcoin] upgrade, which will be explained in more detail in a later subsection.